# Estimation of motion blur kernel parameters using regression convolutional neural networks

### v2

@@ Abstract
Many deblurring and blur kernel estimation methods use MAP or classification deep learning techniques to sharpen an image and predict the blur kernel. We propose a regression approach using neural networks to predict the parameters of linear motion blur kernels. These kernels can be parameterized by its length of blur and the orientation of the blur.This paper will analyze the relationship between length and angle of linear motion blur. This analysis will help establish a foundation to using regression prediction in uniformed motion blur images.
@@ 1 Introduction
Linear motion blur has been studied as a model for camera shake, camera platform movement, and moving objects during imaging. This study was designed as a foundation to model atmospheric turbulence. The effects of atmospheric turbulence are generally modeled by a spatially-varying blurred image which can be modeled as a superposition of linear motion blurs [1, 2]. Linear motion blur kernels are parameterized by length and orientation. In this work, we study the ability to accurately estimate the length and angle parameters of a uniform linear motion blur. The results may then be used as a foundation upon which to build methods to estimate spatially-varying blur parameters and to serve as a baseline in subsequent studies.
@@
$I_{b}=K*I_{s}+N,$ (1)
where $I_{b}$ is the blurred image, $K$ is the blur kernel, $I_{s}$ is the sharp latent image, $N$ is additive noise, and *** is the convolution operator. Note that this formulation assumes a uniform blur since the same kernel $K$ is used across the entire image.
This paper presents an exploration in motion blur kernels including a scope outside of squared odd shaped kernels. The created dataset is used to train VGG16 [3] as a regression network to predict motion blur kernel parameters. The network is analyzed using the $R^{2}$ score which is a common regression analysis as well as a deconvolved error ratio (the error ratio between deconvolving using the predicted vs actual blur kernel).
The organization of this paper is as follows. In Section 2, we discuss previous work, including classical deblurring and deep learning deblurring methods. In Section 3, we describe in detail the exploration of the linear blur kernel parameter space as well as the creation of a blurred dataset for deep learning training. In Section 4, we discuss the process for regression blur parameter prediction and in Section 5, we present results of our regression prediction method and compare them to previous work. Finally, in Section 6, we conclude and briefly discuss our future work.
@@ 2 Related Work
Deblurring and kernel blur estimation has been extensively studied in computer vision, with applications of recovering the sharp image from blurry images caused by camera shake, fast moving objects in frame, or atmospheric turbulence. Prior to deep learning methods, many researchers used a maximum a posteriori (MAP) approach for both blur kernel estimation and deblurring. Two commonly used varieties of MAP are implicit regularization as seen in [4, 5] and explicit edge prediction based approaches like those in [6, 7, 8, 9, 10]. While many approaches have used MAP to estimate both the latent image and the blur kernel, Levin et al. [11, 12] prove that this approach tends to favor the no-blur explanation (i.e., that the kernel is an impulse and the “deblurred” image is the blurry image) instead of converging to the true blur kernel and latent sharp image. Moreover, it is advocated that estimating only the blur kernel is a better approach since there are fewer parameters to estimate than if one were to also estimate the latent sharp image. They use an Expectation-Maximization (EM) framework to optimize their kernel prediction while using either a Gaussian or sparse prior. More recently methods have considered using deep learning approaches to estimate the blur kernel [13], estimate the latent sharp image [14], or both [15].
@@
Edge based approaches in a MAP framework use extracted edges as an image prior in the MAP optimization. Cho et al. [6] introduce a gradient computation using derivatives of the image to compute edges. Money & Kang [8] use shock filters for edge detection. Cho et al. [7] use the Radon transform for edge based analysis; image deblurring may be achieved through the use of a MAP algorithm or by computing the inverse Radon transform informed by the detected edges. Jia [9] uses object boundary transparency as an estimation of edge location. Fergus et al. [10] introduce natural image statistics defined by distributions of image gradients as a prior.
@@
Implicit regularization MAP approaches use different regularization terms to enforce desired image priors. Xu et al. [4] incorporate a regularization term to approximate the $L_{0}$ cost which improves computational speeds over alternative implicit sparse regularizations. This framework alternates between estimating the latent sharp image and the blur kernel in each iteration. Krishnan et al. [5] use a ratio of $L_{1}/L_{2}$ norms as a regularization to estimate the kernel. This helps with the attenuation of high frequencies that blur introduces in an image.
@@
Although many of the previous works discussed above can estimate generalized blur kernels, some work specializes in motion blur kernel prediction. Whyte et al. [16] propose a new method of estimating motion blur by parametrizing a geometric model in terms of rotational velocity of the camera during exposure time. They modify the Fergus et al. [10] algorithm to implement non-uniform deblurring. Hirsch et al. [17] combine projective motion path blur models with efficient filter flow [18] to estimate motion blur at a patch level and deblur by modifying the Krishnan & Fergus [19] algorithm.
@@
Recent methods using deep learning have used various architectures including Convolutional Neural Networks (CNNs) [20, 21, 15, 13], Generative Adversarial Networks (GANs) [14], and Fully Convolutional Networks (FCNs) [22] to tackle the problem of blur kernel prediction and deblurring. Sun et al. [20] use a CNN to classify the best-fit motion blur kernel for each image patch. They use 73 different motion blur kernels and are able to expand up to 361 kernels by rotation of their input image. Gong et al. [22] use an FCN to predict a motion flow map where each pixel has its own motion blur parameters. Li et al. [15] uses a mixture of deep learning and MAP to deblur an image. They use a binary classification CNN which is trained to predict the probability of whether the input is blurry or sharp. The classifier is used as a regularization term of the latent image prior in their MAP framework. Yan & Shao [13] use a two step process to predict blur parameters. Their first step includes a deep neural network to classify the type of blur and their second step is a general regression neural network to predict the parameters of the blur predicted in step one. Gong et al. [22] use an FCN to predict a motion flow map of the blurred image and use non-blind deconvolution to recover the sharp image. While many of these approaches are able to handle spatially varying blur, we focus on a thorough exploration of motion kernel parameters in uniform blur; this will serve as a framework for spatially varying blur in future work.
@@
In this work we focus on motion blur kernels and expand on the complexity of defining a motion blur kernel. We analyze which motion blur kernels exist within different kernel shape sizes rather than using $N\times N$ odd kernels as often implicitly assumed in the implementation, likely for ease of coding. By exploring all the motion blur kernel possibilities for a suite of length and angle combinations, we train a CNN which uses regression prediction instead of classification. We also expand the range of additive noise to analyze Gaussian noise up to a variance of 10%.
@@ 3 Blur Kernels and Blurred Dataset
A linear blur kernel can be parametrized by the length $r$ and orientation $\theta$ of a line. To formulate a discrete blur kernel for application to a digital image using Equation (1), we carefully define the difference between a continuous line and a discrete pixel line. We also explore angle and length combinations that exist within linear motion blur kernels. Finally, we discuss how we use the 2014 COCO dataset to create a new blurred dataset for deep learning training purposes.
@@ 3.1 Continuous and Pixel Lines
In this and following discussions, we use the terminology “line” to refer to a line segment. We use $r$ to denote a Euclidean distance (which could be defined for either a continuous or discrete line), $r_{\infty}$ to denote a Chebyshev distance for a discrete line, $\theta\in(-90,90]$ to denote the orientation (angle) of a continuous line in degrees, and $\phi\in(-90,90]$ to denote the orientation (angle) of a discrete pixel line in degrees. A blur kernel has dimensions $h\times w$ pixels. For a given length $r$, we denote the set of possible continuous line angles as $\Theta\in(-90,90]$ degrees and the set of possible discrete pixel line angles as $\Phi\subseteq\Theta$. A line in a 2-dimensional (2D) continuous domain $\mathbb{R}^{2}$ can be described as a 1-dimensional (1D) object with zero width and length $r$. The orientation of a line can be described by the angle with respect to the horizontal axis; $\theta=0$ thus corresponds to a horizontal line and $\theta=90$ to a vertical line. A pixel line in a 2D discrete domain $\mathbb{Z}^{2}$ is defined on a 2D discrete pixel grid which gives the line a non-zero width and length $r$. Figure 1 illustrates a continuous line with parameters $(r,\theta)=(3,90)$ and a pixel line with parameters $(r,\phi)=(3,90)$.
@@
Figure 1: A continuous line (left) with parameters $(r,\theta)=(3,90^{\circ})$ and pixel line (right) with parameters $(r,\phi)=(3,90^{\circ})$. Note that the pixel line has a width of 1 due to its definition on a 2D discrete grid.
One limitation in defining pixel lines is the interrelation between length and angle of the line. In essence, a shorter pixel line will limit the number of unambiguous angles that a pixel line can have. Figure 2(a) illustrates how multiple continuous lines with angle $\theta\in[0,30]$ can be interpreted by the same horizontal pixel line of Chebyshev length $r_{\infty}=2$. Similarly if we move to a pixel line of Chebyshev length $r_{\infty}=2$ and angle $\phi=45$ (see Figure 2(c)) we can fit many intermediate lines with angles $\theta\in[30,60]$. The four pixel lines illustrated in Figure 2 show the only four pixel lines that can be described for a Chebyshev length $r_{\infty}=2$ pixel line. This means that for smaller length motion blurs, there will be gaps in the angles that can be represented. For example, a length $r_{\infty}=2$ pixel line can only have angles of $\phi\in\{-45,0,45,90\}$ as illustrated in Figure 2.
(a) Two distinct continuous lines in red (with $\theta_{1}=30^{\circ}$ and $\theta_{2}=\phi=0^{\circ}$) described by the same $\phi=0^{\circ}$ pixel line.
(b) Two distinct continuous lines in red (with $\theta_{1}=60^{\circ}$ and $\theta_{2}=\phi=90^{\circ}$) described by the same $\phi=90^{\circ}$ pixel line.
(c) Three distinct continuous lines in red (with $\theta_{1}=30^{\circ}$, $\theta_{3}=\phi=45^{\circ}$, and $\theta_{2}=60^{\circ}$ described by the same $\phi=45^{\circ}$ pixel line.
(d) Three distinct continuous lines in red (with $\theta_{1}=-30^{\circ}$, $\theta_{3}=\phi=-45^{\circ}$, and $\theta_{2}=-60^{\circ}$ described by the same $\phi=45^{\circ}$ pixel line.
@@
Figure 2: Multiple continuous lines can be represented by the same pixel line of Chebyshev length $r_{\infty}=2$. Example continuous lines are shown in red, overlaid on the pixel line kernel that describes those continuous lines.
@@
We also note that blur kernels may have two additional characteristics different from common conventions in image filter kernels: they may be non-square in dimension and they may be even in one or more dimensions. The $\phi=0$ and $\phi=90$ kernels in Figure 2 are $1\times 2$ and $2\times 1$ pixels, respectively; these two kernels are non-square and even in one dimension. The $\phi=\pm 45$ kernels in Figure 2 are both $2\times 2$ pixels; these two kernels are square but even in both dimensions. The assumption of square or odd-sized convolution kernels in the blur model described in Equation (1) is not mathematically necessary. The preference for square kernels is related to the symmetry of features both horizontally and vertically. The preference for odd-sized kernels is that there is an unambiguous center point to the kernel. This center point is commonly assumed to be associated with the pixel that is being processed.
@@ 3.2 Blur Kernel Creation
We explore the gaps in angles resulting from representation of continuous lines as pixel lines. We study a range of Euclidean lengths $r=2,3,\ldots,100$ and angles $\theta=-89,-88,\ldots,90$, and gather the resulting and unique $(r,\phi)$ discrete line parameter pairs.
We calculate the shape of an $h\times w$ kernel corresponding to an $(r,\theta)$ continuous line as follows:
$h=\begin{cases}\lceil r\cos(\theta)\rceil,&\cos(\theta)\neq 0\\ 1,&\cos(\theta)=0\\ \end{cases}$ (2)
$w=\begin{cases}\lceil r\sin(\theta)\rceil,&\sin(\theta)\neq 0\\ 1,&\sin(\theta)=0,\\ \end{cases}$ (3)
@@
For each length $r=2,3,\ldots,100$, we generate a discrete blur kernel associated with a continuous line of length $r$ and angle $\theta=0,1,\ldots,90$. We then group all the angles $\theta\in\Theta$ that resulted in the same pixel line (e.g., see Figure 2). We must define a unique single discrete angle $\phi$ for each group of angles $\Theta$ for use as a training label:
$\phi=\begin{cases}0,&0\in\Theta\\ 90,&90\in\Theta\\ \lceil\text{median}(\Theta)\rceil,&else\end{cases}.$ (4)
Note that this definition of $\phi$ has two special cases. If the collection contains $\theta=0$, i.e., $0\in\Theta$, we know that the kernel is consistent with a horizontal line (see Figure 2(a)) and we assign $\phi=0$. Similarly, if the collection contains $\theta=90$, i.e., $90\in\Theta$, we know that the kernel is consistent with a vertical line (see Figure 2(b)) and we assign $\phi=90$. These special cases avoid the median operator assigning an erroneous label to a horizontal or vertical line.
We compute all possible $(r,\phi)$ combinations by generating the blur kernel for $r=2,3,\ldots,100$ and $\theta=0,1,\ldots 90$ and computing the resulting blur kernel angle $\phi$ using Equation (4). We span only positive angles $\theta\in[0,90]$ since we know that the negative kernels will be symmetric versions of the positive kernels (see also Figure 2(c) and (d)). The resulting combinations of $(r,\phi)$ parameters are the parameters of the unique discrete pixel lines. These parameter values are also what will be used as labels to train a network to predict the blur parameters from a blurry image (Section 4), using $(r,\phi)=(1,0)$ as labels for a non-blurred image.
(a) Unique combinations $(r,\phi)$ for discrete pixel lines.
(b) Number of unique angles $\phi$ for each length $r$ in a discrete pixel line.
Figure 3: Length and angle pairs $(r,\phi)$ that are describable by a discrete pixel line. (a) Length $r$ and angle $\phi$ for all possible discrete blur kernels for $r\in[2,100]$ and $\phi\in(-90,90]$. (b) Number of angles that can be represented for a given length kernel. Note that shorter length kernels have many fewer possible angles that they can represent, including the four possible angles of $-45$, 00, $45$, and $90$ for $r=2$.
Figure 3(a) illustrates the $(r,\phi)$ combinations as a scatter plot, where we note the existence of gaps where a continuous line cannot be described by a pixel line. As expected those gaps are larger for smaller lengths, indicating that there are limited unique pixel lines (blur kernels) for shorter blur lengths. Figure 3(b) plots the number of unique angles $\phi$ versus length $r$ where we note that we must have a line of approximately 70 pixels or larger in order to represent all orientations $\phi=-89,-88,\ldots,90$. We denote the unique combinations as $(r_{u},\phi_{u})$ and use these values to create a blurred dataset.

### v3

@@ Abstract
Many deblurring and blur kernel estimation methods use a maximum a posteriori (MAP) approach or deep learning-based classification techniques to sharpen an image and/or predict the blur kernel. We propose a regression approach using convolutional neural networks (CNNs) to predict parameters of linear motion blur kernels, the length and orientation of the blur. We analyze the relationship between length and angle of linear motion blur that can be represented as digital filter kernels. A large dataset of blurred images is generated using a suite of blur kernels and used to train a regression CNN for prediction of length and angle of the motion blur. The coefficients of determination for estimation of length and angle are found to be greater than or equal to 0.89, even under the presence of significant additive Gaussian noise, up to a variance of 10% (SNR of 10 dB). Using our estimated kernel in a non-blind image deblurring method, the sum of squared differences error ratio demonstrates higher cumulative histogram values than comparison methods, with most test images yielding an error ratio of less than or equal to 1.25.
11footnotetext: The official version of this paper appears as Luis G. Varela, Laura E. Boucheron, Steven Sandoval, David Voelz, Abu Bucker Siddik, “Estimation of motion blur kernel parameters using regression convolutional neural networks,” J. Electron. Imaging 33(2), 023062 (2024); doi: 10.1117/1.JEI.33.2.023062.22footnotetext: Copyright 2024 Society of Photo-Optical Instrumentation Engineers. One print or electronic copy may be made for personal use only. Systematic reproduction and distribution, duplication of any material in this paper for a fee or for commercial purposes, or modification of the content of the paper are prohibited.
@@ 1 Introduction
Linear motion blur has been studied as a model for camera shake, camera platform movement, and moving objects during imaging [1, 2]. The studies presented within were designed as an initial exploration of a regression convolutional neural network (CNN) to estimate linear motion blur parameters from blurry images. The effects of atmospheric turbulence in an image are generally represented by a spatially-varying blur [3, 4], while recent work has demonstrated that spatially-varying blurred images can be modeled as a superposition of locally linear motion blurs [5]. One way to parametrize linear motion blur kernels is by length and orientation. In this work, we study the ability to accurately estimate the length and angle parameters of a uniform linear motion blurred image. The results may then be used as a foundation upon which to build methods to estimate spatially-varying blur parameters and to serve as a baseline in subsequent studies.
@@
$I_{\text{b}}=K*I_{\text{s}}+N,$ (1)
where $I_{\text{b}}$ is the blurred image, $K$ is the blur kernel, $I_{\text{s}}$ is the sharp latent image, $N$ is additive noise, and * is the convolution operator. This formulation assumes a uniform blur because the same kernel $K$ is applied across the entire image.
The main contributions of this paper are a thorough exploration of linear motion blur kernels and development of a regression CNN for blind prediction of motion blur parameters from blurry images. We study the space of motion blur parametrized by length and angle, specifically the subset of linear motion blur that can be described as 2D discrete motion blur kernels. Furthermore, we include motion blur kernels that are not described as square odd-sized kernels and study complications that arise in defining 2D discrete motion blur kernels. To that end we create a dataset of blurred images spanning a large range of possible blur kernels for use in training and testing a regression CNN based on the VGG16 [6] network to predict linear motion blur kernel parameters. The network is analyzed using the coefficient of determination $R^{2}$ score, which is a common metric to quantify performance of regression analysis, as well as a deconvolved error ratio (the ratio of the error when deconvolving using the predicted blur kernel to the error when deconvolving using the actual blur kernel) [7].
The organization of this paper is as follows. Section 2 gives a discussion on previous work including classical deblurring and deep learning deblurring methods. Section 3 contains a detailed description of the exploration of the linear blur kernel parameter space as well as the creation of a blurred dataset for deep learning training. Section 4 presents the proposed regression CNN for blur parameter prediction and Sec. 5 presents results of our regression prediction method along with comparison to previous work. Finally, in Sec. 6 we provide a conclusion and briefly discuss our future work.
@@ 2 Related Work
Deblurring and kernel blur estimation have been extensively studied in computer vision with applications of recovering the sharp image from blurry images caused by camera shake, fast moving objects in frame, or atmospheric turbulence. Prior to deep learning methods, many researchers used a maximum a posteriori (MAP) approach for both blur kernel estimation and deblurring. Two commonly used varieties of MAP are implicit regularization as seen in [8, 9] and explicit edge prediction based approaches like those in [10, 11, 12, 13, 1]. While many approaches have used MAP to estimate both the latent image and the blur kernel, Levin et al. [14, 7] prove that this approach tends to favor the no-blur explanation (i.e., that the kernel is an impulse and the “deblurred” image is the blurry image) instead of converging to the true blur kernel and latent sharp image. Moreover, it is advocated in [14, 7] that estimating only the blur kernel is a better approach since there are fewer parameters to determine than if one were to also estimate the latent sharp image. Levin et al. [14, 7] use an expectation-maximization (EM) framework to optimize kernel prediction with either a Gaussian or sparse prior. More recently methods have considered deep learning approaches to estimate the blur kernel [15], estimate the latent sharp image [16], or both [17].
@@
In the first variety of MAP-based methods, edge-based approaches in a MAP framework use extracted edges as an image prior in the MAP optimization. Cho et al. [10] introduced a gradient computation using derivatives of the image to compute edges. Money & Kang [12] used shock filters for edge detection. Cho et al. [11] used the Radon transform for edge based analysis; image deblurring may use a MAP algorithm or the inverse Radon transform informed by the detected edges. Jia [13] used object boundary transparency as an estimation of edge location. Fergus et al. [1] introduced natural image statistics defined by distributions of image gradients as a prior.
@@
In the second variety of MAP-based methods, implicit regularization MAP approaches use different regularization terms to enforce desired image priors. Xu et al. [8] incorporated a regularization term to approximate the $L_{0}$ cost which improves computational speeds over alternative implicit sparse regularizations. The framework in [8] alternated between estimating the latent sharp image and the blur kernel in each iteration. Krishnan et al. [9] used a ratio of $L_{1}/L_{2}$ norms as a regularization to estimate the kernel. This helps with the attenuation of high frequencies that blur introduces in an image.
@@
Although many of the previous works discussed above involve the estimation of generalized blur kernels, some work specializes in motion blur kernel prediction. Whyte et al. [18] proposed a new method to estimate motion blur by parametrizing a geometric model in terms of rotational velocity of the camera during exposure time. They modified the Fergus et al. [1] algorithm to implement non-uniform deblurring. Hirsch et al. [19] combined projective motion path blur models with efficient filter flow [20] to estimate motion blur at a patch level and deblur by modifying the Krishnan & Fergus [21] algorithm.
@@
Recent methods using deep learning have used various architectures including convolutional neural networks (CNNs) [22, 23, 17, 15, 24], encoder-decoder networks [25], generative adversarial networks (GANs) [16], and fully convolutional networks (FCNs) [26] to tackle the problems of blur kernel prediction and deblurring. Sun et al. [22] used a CNN to classify the best-fit motion blur kernel for each image patch. They used 73 different motion blur kernels and were able to expand to 361 kernels by rotation of their input image. Xu et al. [23] used a CNN to recover sharp edges from blurred images and then estimated the blur kernel from the blurry and recovered edges via a MAP optimization. Li et al. [17] used a mixture of deep learning and MAP estimation to deblur an image, using a binary classification CNN trained to predict the probability of whether the input is blurry or sharp. The classifier is used as a regularization term of the latent image prior in a MAP framework to deblur. Yan & Shao [15] used a two step process to predict blur parameters: first, a deep neural network classified the type of blur and second, a general regression neural network predicted the parameters of the blur. Nasonov and Nasonova [24] used a CNN to predict length and angle for linear motion blur with several formulations of the output values, although they noted difficulty in the simultaneous estimation of length and angle. Carbajal et al. [25] used an encoder and two decoders to predict motion kernel bases for the image and mixing coefficients for each pixel. Zhang et al. [16] used a conditional GAN to deblur images, combining an adversarial loss and a perceptually-motivated content loss. Gong et al. [26] used an FCN to predict a motion flow map where each pixel has its own motion vector estimate. While many of these approaches are able to handle spatially varying blur, we focus on a thorough exploration of motion kernel parameters in uniform blur; this will serve as a framework for spatially varying blur in future work.
@@
There are three main contributions to this work. First, we study complications that arise in defining a motion blur kernel and, in particular, we analyze which motion blur kernels exist within different kernel shapes rather than using square odd-sized kernels as often implicitly assumed in the implementation—likely for ease of coding. Second, by exploring all motion blur kernel possibilities for a suite of length and angle combinations, we train a CNN which uses regression prediction instead of classification. By formulating the prediction as a regression, we train two output nodes (one for length and one for angle) which can span the full range of possible parameter values and can be trained on any granularity of parameters. Use of a classification network would necessitate 13,034 output nodes for the granularity considered here (99 lengths and 180 angles) and modification of the network architecture to train on any other granularity or post-hoc computation (e.g., as in [22]) to define predictions at another granularity. Use of regression also may alleviate some of the issues in simultaneous prediction of length and angle noted in [24]. Third, we expand the range of additive noise beyond that studied for other methods [14, 11, 23, 17, 9, 18] to analyze Gaussian noise up to a variance of 10% (10 dB SNR).
@@ 3 Blur Kernels and Blurred Dataset
A linear blur kernel can be parametrized by the length and orientation angle of a line. To formulate a discrete blur kernel for application to a digital image using Eq. (1), we specifically distinguish between a continuous line and a discrete pixel line. Next, we explore angle and length combinations that exist within linear motion blur kernels. Finally, we describe how the 2014 COCO dataset [27] is used to create a new blurred dataset for deep learning training and testing.
@@ 3.1 Continuous and Pixel Lines
Here, the term “line” refers to a line segment, $r$ denotes a Euclidean distance (or a continuous or discrete line), $r_{\infty}$ is a Chebyshev distance for a discrete line, $\theta\in(-90,90]$ denotes the orientation (angle) of a continuous line with respect to the horizontal axis in units of degrees, and $\phi\in(-90,90]$ is the angle of a discrete pixel line with the same conventions. For a given length $r$, the set of possible continuous line angles is denoted $\Theta\in(-90,90]$ and the set of possible discrete pixel line angles is $\Phi\subseteq\Theta$. A line in a 2D continuous domain $\mathbb{R}^{2}$ is a 1D object with zero width and length $r$. On the other hand, a pixel line in a 2D discrete domain $\mathbb{Z}^{2}$ is defined on a discrete pixel grid which gives the line a non-zero width and length $r$. Figure 1 shows a continuous line with parameters $(r,\theta)=(3,90)$ and a pixel line with parameters $(r,\phi)=(3,90)$.
@@
Figure 1: (a) A continuous line with parameters $(r,\theta)=(3,90^{\circ})$ and (b) pixel line with parameters $(r,\phi)=(3,90^{\circ})$. Note that the pixel line has a width of 1 due to its definition on a 2D discrete grid.
@@
Figure 2: Multiple continuous lines can be represented by the same pixel line of Chebyshev length $r_{\infty}=2$. Example continuous lines are shown in red, overlaid on the pixel line kernel (the squares) that describes those continuous lines, where white squares (pixels) correspond to the line and black squares (pixels) correspond to the absence of a line. (a) Two continuous lines (with $\theta_{1}=30^{\circ}$ and $\theta_{2}=\phi=0^{\circ}$) described by the same $\phi=0^{\circ}$ pixel line. (b) Two continuous lines (with $\theta_{1}=60^{\circ}$ and $\theta_{2}=\phi=90^{\circ}$) described by the same $\phi=90^{\circ}$ pixel line. (c) Three continuous lines (with $\theta_{1}=31^{\circ}$, $\theta_{3}=\phi=45^{\circ}$, and $\theta_{2}=59^{\circ}$ described by the same $\phi=45^{\circ}$ pixel line. (d) Three continuous lines (with $\theta_{1}=-31^{\circ}$, $\theta_{3}=\phi=-45^{\circ}$, and $\theta_{2}=-59^{\circ}$ described by the same $\phi=-45^{\circ}$ pixel line.
One limitation in defining pixel lines is the interrelation between length and angle. In essence, a shorter pixel line will have a limited number of unambiguous angles. Figure 2(a) illustrates how multiple continuous lines with angle $\theta\in[0,30]$ can be interpreted by the same horizontal pixel line of length $r_{\infty}=2$. Similarly, a pixel line of length $r_{\infty}=2$ and angle $\phi=45$ [Fig. 2(c)] can describe many intermediate lines with $\theta\in(30,60)$. The four pixel lines in Fig. 2 show the only pixel lines that can be described for length $r_{\infty}=2$. This means that, for shorter length motion blurs, there will be gaps in the angles that can be represented. For example, a length $r_{\infty}=2$ pixel line can only have angles $\phi\in\{-45,0,45,90\}$ as shown in Fig. 2.
@@
Conventional convolutional kernels are typically assumed to be square and odd in dimension. The preference for square kernels is related to the horizontal and vertical symmetry of features. The preference for odd-sized kernels is to have an unambiguous center point, as the center point is commonly assumed to be associated with the pixel being processed. The assumption of square or odd-sized kernels in the blur model of Eq. (1) is not mathematically necessary, however, and we consider blur kernels that may be non-square and/or even in one or more dimension.
@@ 3.2 Blur Kernel Creation
Figure 3: (a) Unique length and angle combinations $(r,\phi)_{\text{u}}$ for all possible discrete blur kernels for $r\in[2,100]$ and $\phi\in(-90,90]$. Shorter length kernels have many fewer possible angles that they can represent, including the four possible angles of $\phi\in\{-45,0,45,90\}$ for $r=2$. (b) Number of angles that can be represented for a given kernel length. Shorter length kernels have many fewer possible angles that they can represent.
To explore the gaps in angles resulting from representation of continuous lines as pixel lines, we assume a range of Euclidean lengths $r=2,3,\ldots,100$ and angles $\theta=-89,-88,\ldots,90$, and gather the resulting unique $(r,\phi)$ discrete line parameter pairs, illustrated in Fig. 3. We assume the $h\times w$ kernel corresponding to an $(r,\theta)$ continuous line has $h$ and $w$ defined as
$h=\begin{cases}\lceil r\cos(\theta)\rceil,&\cos(\theta)\neq 0\\ 1,&\cos(\theta)=0\\ \end{cases},$ (2)
$w=\begin{cases}\lceil r\sin(\theta)\rceil,&\sin(\theta)\neq 0\\ 1,&\sin(\theta)=0\\ \end{cases},$ (3)
@@
For each length $r=2,3,\ldots,100$, we generate a discrete blur kernel associated with a continuous line of length $r$ and angle $\theta=0,1,\ldots,90$. Noting that multiple angles $\theta$ may result in the same discrete blur kernel (see Fig. 2), a single discrete angle $\phi$ must be defined for each unique discrete blur kernel for use as a training label. We thus quantize the continuous angle $\theta$ for a given blur kernel as
$\phi=\begin{cases}0,&0\in\Theta\\ 90,&90\in\Theta\\ \lceil\text{median}(\Theta)\rceil,&\text{else}\end{cases},$ (4)
[…]
We compute all possible $(r,\phi)$ combinations by generating the blur kernel for $r=2,3,\ldots,100$ and $\theta=0,1,\ldots,90$ and computing the resulting blur kernel angle $\phi$ using Eq. (4). We span only positive angles $\theta\in[0,90]$ since kernels with negative angles will be symmetric versions of the kernels with positive angles. The resulting values $r$ and $\phi$ that yield unique discrete pixel lines are denoted as $(r,\phi)_{\text{u}}$. These parameter values are also used as labels to train a network to predict the blur parameters from a blurry image (Sec. 4) using $(r,\phi)=(1,0)$ as labels for a non-blurred image.
Figure 3 illustrates the $(r,\phi)_{\text{u}}$ combinations as a scatter plot, where we note the existence of gaps where a continuous line cannot be described by a pixel line. As expected, gaps are larger for smaller lengths, indicating limited unique pixel lines (blur kernels) for shorter blur lengths. Figure 3 plots the number of unique angles $\phi$ versus length $r$ where we note that a line must have approximately 70 pixels or more in order to represent all orientations $\phi=-89,-88,\ldots,90$. This exploration of the $(r,\phi)$ space explored 17,820 possible $(r,\theta)$ lines resulting in 13,034 unique $(r_{,}\phi)_{\text{u}}$ kernels; this is orders of magnitude larger than other explorations, e.g., the 361 kernels in [22].